OpenStudy (anonymous):

A Ferris Wheel has a diameter of 56 metres, and one revolution takes 2.5 minutes to complete. Riders may see Niagara Falls if they are highter than 50 metres above the ground. a) Sketch and label a cycle of the graph that represents the height of the rider above the ground, as a function of time, assuming that the rider gets on at a height of 0.5 metres at the bottom of the Ferris wheel at t=0 minutes. b) Determine the equation of the function, using h for height and t for time. c) Are riders able to see the Falls after one minute on the Ferris wheel? Show your solution?

OpenStudy (anonymous):

lord i can see why no one wants to do this, but i think we can do it

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

first of it is clearly a trig function right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

translated up

OpenStudy (anonymous):

you know the period is 2.5 and you know the diameter is 56 so the amplitude will be 28

OpenStudy (anonymous):

for example \[28\cos(x)\] goes from 28 to -28

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

continue...

OpenStudy (anonymous):

you are going to have to use \[-\cos(x)\] because you want the minimum at t = 0 where cosine is -1

OpenStudy (anonymous):

so it is going to look like \[-28\cos(bx)+c\]

OpenStudy (anonymous):

so that is equation?

OpenStudy (anonymous):

i would have said \[c=28\] but that doesn't work because you are .5 up to start, so we can use \[-28\cos(bx)+28.5\] to insure that we start at .5

OpenStudy (anonymous):

then we have to find b

OpenStudy (anonymous):

and we know the period is 2.5 minutes, and the formula for the period is \[\frac{2\pi}{b}=P\]so \[\frac{2\pi}{b}=2.5\] solve for b and get \[b=\frac{4\pi}{5}\] check my algebra

OpenStudy (anonymous):

so i believe your function should be... \[-28\cos(\frac{4\pi}{5}x)+28.5\]

OpenStudy (anonymous):

how are we doing so far?

OpenStudy (anonymous):

yeah i got some ? i trying to understand i think u r right thx

OpenStudy (anonymous):

not done yet though

OpenStudy (anonymous):

can u solve my next question?

OpenStudy (anonymous):

here is a nice graph to show what we got. it looks right http://www.wolframalpha.com/input/?i=-28cos%284pi+x%2F5%29+%2B+28.5

OpenStudy (anonymous):

this graph don't have cycle.

OpenStudy (anonymous):

now we have to see how high they are at t = 1, so put t = 1 and find the hight. oh it says write height as a function of t, so you should write \[h(t)=-28\cos(\frac{4\pi}{5}t)+28.5\]

OpenStudy (anonymous):

yes the graph has a cycle. it goes up and down and up and down...

OpenStudy (anonymous):

now if we replace t by 1 we get \[h(1)=-28\cos(\frac{4\pi}{5})+28.5=51.1525...\] from a calculator, so yes, they can see niagra falls http://www.youtube.com/watch?v=9pQii1L8fGk&feature=related

OpenStudy (anonymous):

i got it